The decomposition of ammonium hydrogen sulfide
NH4HS (s) <=> NH3 (g) + H2S (g)
is an endothermic process. A 6.1589-g sample of the solid is placed in an evacuated 4.000-L vessel at exactly 24°C. After equilibrium has been established, the total pressure inside is 0.709 atm. Some solid NH4HS remains in the vessel.
(a) What is the KP for the reaction?
(b) What percentage of the solid has decomposed?
1
Expert's answer
2020-11-17T10:32:55-0500

a) Pressure at equilibrium = 0.709atm

Ptotal = PNH3 + PH2S

0.709 = x+x

0.709= 2x

x= 0.355atm

Kp= (x)(x)= 0.5552 = 0.1260


b) PV = nRT

n = PV/RT = (0.355 × 4 )/ (0.0821× 297)

= 0.0582mol

Mass of decomposed solid = mass of NH3 and H2S at equilibrium

Molar mass of NH3 and H2S is 17.031 and 34.081 respectively

Total mass of NH3 and H2S = 0.0582(17.031) + 34.081(0.0582)= 2.975g

% decomposed = 2.975/ 6.1589 × 100% = 48.3%